find the greatest common Divison (GCD)36,48, and 84
Answer by stanbon(75887)

find the greatest common Divison (GCD)
---
36 = 2^2*3^2
------
48 = 2^4*3
------
84 = 2^2*3*7
====================
GCD = 2^2*3 = 12
=================
Cheers,
Stan H.